While walking on one bank of a 25 m wide canal you see a child 50m down the canal on the far bank slip, hit his head, and fall into the water. If you can swim 1.5m/s through the water and can run 5.0m/s along the side of the pool, what distance should you swim through the water to minimize the time to get to the child?


1
Expert's answer
2021-03-26T10:28:54-0400

The length is 50 m and the width is 25 m.




Distance AB


"AB^2=50^2+25^2"

Let I run "x\ m" along the side of the pool. Then I swim


"BD=\sqrt{(50-x)^2+25^2}"

the time to get to the child is


"t=t(x)=\dfrac{x}{5}+\dfrac{\sqrt{(50-x)^2+25^2}}{1.5}, 0\leq x\leq 50"

Find the first derivative with respect to "x"


"t'(x)=\dfrac{1}{5}-\dfrac{50-x}{1.5\sqrt{(50-x)^2+25^2}}"

Find the critical number(s)


"t'(x)=0=>\dfrac{1}{5}-\dfrac{50-x}{1.5\sqrt{(50-x)^2+25^2}}=0"

"10(50-x)=3\sqrt{(50-x)^2+25^2}"

"100(50-x)^2=9(50-x)^2+9(625)"

"91(50-x)^2=9(625)"

"x=50-\dfrac{75}{\sqrt{91}}"

First Derivative Test

If "0<x<50-\dfrac{75}{\sqrt{91}}, t'(x)<0, t(x)" decreases

If "50-\dfrac{75}{\sqrt{91}}<x<50, t'(x)>0,t(x)" increases.

Then function "t(x)" has a local minimum at "x=50-\dfrac{75}{\sqrt{91}}."

Sincd the function "t(x)" has the only extremum, then the function "x(t)" has the absolute minimum at "x=50-\dfrac{75}{\sqrt{91}}."

"BD=\sqrt{(50-50+\dfrac{75}{\sqrt{91}})^2+25^2}=\dfrac{250}{\sqrt{91}}"

"\approx26.207(m)"

I should swim 26.207 m through the water to minimize the time to get to the child.